a match box 4 cm long,2.5 cm wide and 15 cm high is to be covered exactly with coloured paper how much paper will be required?
Answer by JBarnum(2146)

First off that's one wired looking box of matches...lol
It's best to draw a picture diagram to help you see how many sides there are and which measurements are on each side
You should have: two (4 x 2.5) , two (4 x 15) , and two (15 x 2.5) sides
Finding the sum of all the areas of each side will give you total surface area
Your equation should look like:
2(4*2.5)+2(4*15)+2(15*2.5)=S
2(10)+2(60)+2(37.5)=S
20+120+75=S
215=S
Total surface area is 215cm